# Deep Spectral Prior

### v1

@@ 2.2 Spectral Regularisation and Implicit Bias of Deep Spectral Prior
Moreover, if the network $f_{\theta}$ exhibits spectral bias, then the components of $\mathcal{F}(u_{t})$ at low spatial frequencies converge faster than those at high frequencies. (See proof in Appendix.)
2.3 Optimisation Properties of Deep Spectral Prior
Having shown that DSP introduces a desirable regularisation mechanism, we now examine its optimisation properties. Specifically, we verify that the frequency loss is smooth, locally stable, and conducive to convergence under standard first-order methods. This theoretical part explains why DSP is compatible with common optimisers like Adam and why its convergence dynamics remain stable across diverse inverse tasks.
Theorem 2.2 (Smoothness and Local Stability of our Frequency Loss).
Let $f_{\theta}:\mathbb{R}^{d}\to\mathbb{R}^{n}$ be a neural network that is continuously differentiable with respect to parameters $\theta\in\mathbb{R}^{p}$. Let $\mathcal{A}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ be a fixed linear corruption operator, and let $\mathcal{F}:\mathbb{R}^{m}\to\mathbb{C}^{m}$ denote the discrete Fourier transform (DFT). For a fixed input $z\in\mathbb{R}^{d}$ and observed corrupted image $y\in\mathbb{R}^{m}$, define the DSP frequency loss as: $\mathcal{L}_{\mathrm{DSP}}(\theta):=\left\|\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(y)\right\|_{2}^{2}.$ It follows that:
  1. 1.
    $\mathcal{L}_{\mathrm{freq}}$ is continuously differentiable with respect to $\theta$;
  2. 2.
    $\mathcal{L}_{\mathrm{freq}}$ is locally Lipschitz in $\theta$;
  3. 3.
    Any local minimum $\theta^{*}$ satisfies the first-order optimality condition $\nabla_{\theta}\mathcal{L}_{\mathrm{freq}}(\theta^{*})=0$;
  4. 4.
    In a neighborhood of any stationary point $\theta^{*}$, gradient-based optimisation (e.g., gradient descent or Adam with sufficiently small step size) yields descent in the frequency domain,
    $\mathcal{L}_{\mathrm{freq}}(\theta_{t+1})\leq\mathcal{L}_{\mathrm{freq}}(\theta_{t}),$
    for all $t$ sufficiently large.                                                   (See proof in Appendix.)
To solidify the theoretical foundation of DSP, it is critical to understand how frequency-domain objectives relate to their spatial-domain counterparts. The following result establishes that, under the unitary Fourier transform, minimising the DSP loss with complex-valued frequency alignment is strictly equivalent to minimising the traditional pixel-wise loss. This equivalence is not merely cosmetic: it guarantees that the shift to the frequency domain does not alter the optimisation landscape, but rather recasts it in a basis aligned with the spectral properties of images. In doing so, DSP inherits the guarantees of pixel-space reconstruction while gaining interpretability, robustness, and inductive alignment with convolutional architectures.
Let $x,\hat{x}\in\mathbb{R}^{m}$, and let $\mathcal{F}:\mathbb{R}^{m}\rightarrow\mathbb{C}^{m}$ denote the unitary discrete Fourier transform. Then:
$\|x-\hat{x}\|_{2}^{2}=\|\mathcal{F}(x)-\mathcal{F}(\hat{x})\|_{2}^{2}.$
In particular, for any neural network output $f_{\theta}(z)$ and measurement $y$, the complex-valued frequency-domain loss $\mathcal{L}_{\mathrm{DSP}}(\theta):=\left\|\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(y)\right\|_{2}^{2}$ is exactly equal to the pixel-space reconstruction error $\|f_{\theta}(z)-y\|_{2}^{2}$. Thus, minimising $\mathcal{L}_{\mathrm{freq}}$ guarantees pixel-accurate recovery. (See proof in Appendix.)
These properties guarantee that DSP remains compatible with standard deep learning practices and optimisers, while extending the functional landscape with frequency-aligned descent.
2.4 Spectral Geometry and Signal Structure
To further understand the effect of DSP on reconstruction quality, we analyse the bias–variance tradeoff in the frequency domain. This classical decomposition provides a quantitative lens on how DSP balances signal fidelity and robustness. In particular, we are interested in how frequency alignment affects the distribution of bias and variance across the spectrum. The following proposition formalises this decomposition, revealing how spectral bias in convolutional networks interacts with the DSP objective.
Proposition 2.2 (Frequency-Domain Bias–Variance Tradeoff).
Let $\hat{x}_{\theta}$ be the network output trained to minimise the DSP loss $\mathcal{L}(\theta):=\left\|\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(y)\right\|_{2}^{2},$ Then the expected error at convergence in the Fourier domain decomposes as
$\mathbb{E}\left[\left\|\mathcal{F}(\hat{x}_{\theta})-\mathcal{F}(x)\right\|_{2}^{2}\right]=\underbrace{\left\|\mathbb{E}\left[\mathcal{F}(\hat{x}_{\theta})\right]-\mathcal{F}(x)\right\|_{2}^{2}}_{\text{Bias}^{2}}+\underbrace{\mathbb{E}\left[\left\|\mathcal{F}(\hat{x}_{\theta})-\mathbb{E}[\mathcal{F}(\hat{x}_{\theta})]\right\|_{2}^{2}\right]}_{\text{Variance}}.$
Due to the spectral bias of convolutional networks, the bias term decays faster at low frequencies, while the variance remains higher in high-frequency components—resulting in implicit regularisation through spectral ordering. (See proof in Appendix.)
This result clarifies the regularisation mechanism underlying DSP. Due to the spectral bias of neural networks, low-frequency components are fitted more consistently across different training runs, resulting in lower variance and faster convergence. Conversely, high-frequency components exhibit greater variance and are less stable, especially in the presence of noise. Crucially, the frequency-domain loss implicitly reflects this structure: although the loss penalises all frequency components equally in form, the network’s inductive bias causes the optimisation to prioritise smoother, low-frequency solutions. As a result, DSP achieves a favourable bias–variance profile without requiring explicit filtering or handcrafted priors. This behaviour contrasts sharply with classical DIP, where the pixel-space loss treats all spatial details equally, often leading to overfitting unless training is manually stopped. DSP, in contrast, exhibits natural spectral regularisation driven by both the loss and the optimiser’s dynamics.
While the previous bias–variance analysis quantifies how DSP balances expressivity and stability across frequency bands, we now shift to a geometric interpretation. Specifically, we ask: what set of signals does DSP consider valid reconstructions? The frequency-domain loss defines an implicit constraint: reconstructed images must match the observed measurement in the Fourier domain. This constraint naturally induces a manifold of frequency-consistent candidates. The following proposition formalises this idea, showing that DSP can be viewed as projecting the network output onto this manifold—thus ensuring spectral fidelity without requiring explicit supervision or regularisation.
Proposition 2.3 (Implicit Projection onto the Frequency Consistency Manifold).
Let $f_{\theta}(z)\in\mathbb{R}^{n}$ denote the network output, and let $\mathcal{A}\in\mathbb{R}^{m\times n}$ be a linear degradation operator. Define the frequency consistency manifold:
$\mathcal{M}_{y}:=\left\{x\in\mathbb{R}^{n}:\mathcal{F}(\mathcal{A}x)=\mathcal{F}(y)\right\}.$
Then, at convergence, the DSP optimisation seeks a point $\hat{x}_{\theta}$ satisfying:
$\hat{x}_{\theta}\approx\arg\min_{x\in\mathcal{M}_{y}}\|x-f_{\theta}(z_{0})\|^{2},$
where $z_{0}$ is the fixed network input. That is, DSP approximates a projection of the unconstrained network output onto the frequency-consistent manifold $\mathcal{M}_{y}$. (See proof in Appendix.)
This proposition offers a powerful geometric lens through which to interpret DSP. Rather than merely fitting data in an unconstrained optimisation landscape, DSP implicitly restricts solutions to lie on a low-dimensional manifold defined by spectral consistency with the measurement. This mechanism ensures that reconstructions preserve the frequency characteristics of the observed data while filtering out artefacts that violate this constraint. Importantly, this projection is implicit—it arises naturally from the optimisation dynamics, without requiring handcrafted priors or external supervision. By aligning the network’s output with a well-structured manifold, DSP stabilises the reconstruction process and improves generalisation across inverse problems. This interpretation also highlights how DSP turns frequency alignment into a form of data-driven constraint enforcement, distinguishing it fundamentally from pixel-space DIP which lacks such an explicit geometric prior.
A critical vulnerability of classical DIP is its tendency to overfit noise, especially when the corruption involves high-frequency perturbations. This is a direct consequence of its pixel-space loss, which treats all frequencies uniformly. In contrast, DSP exhibits a form of inherent robustness to such noise—without requiring explicit denoising modules or handcrafted priors. The next formalise this observation: under the spectral bias of deep convolutional networks, DSP naturally filters out high-frequency components of the noise and prioritises signal recovery in low-frequency bands. This explains why DSP maintains reconstruction quality even in challenging noisy regimes.
Proposition 2.4 (Stability to High-Frequency Noise).
Let $y=\mathcal{A}x+\eta$, where $\eta\in\mathbb{R}^{m}$ is additive noise. Let $\mathcal{F}$ denote the discrete Fourier transform, and suppose $\eta$ is supported primarily in high-frequency bands. Consider the DSP loss: $\mathcal{L}(\theta):=\left\|\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(y)\right\|_{2}^{2}.$ Then, under spectral bias of $f_{\theta}$, the optimisation trajectory satisfies:
$\left\|\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(x)\right\|_{2}^{2}\leq\left\|\mathcal{F}(\eta_{<k})\right\|_{2}^{2}+\mathcal{O}(t^{-\alpha}),$
where $\eta_{<k}$ is the projection of $\eta$ onto low-frequency components, $t$ is the iteration index, and $\alpha>0$ is a constant determined by network architecture and depth. (See proof in Appendix.)
This result confirms that DSP acts as a spectral filter during optimisation, concentrating its fitting capacity on the most informative frequency bands. Since high-frequency noise typically dominates the tail of the spectrum, its effect on the loss diminishes over time. The bound demonstrates that reconstruction error is ultimately governed by the low-frequency content of the noise, which tends to be minimal in natural settings. Practically, this means DSP avoids the need for early stopping, denoising pre-processing, or post-hoc regularisation. The network’s own training dynamics, when using the frequency loss, yield an implicit denoising mechanism.
3 Experimental Results
We evaluate DSP on blind denoising, restoration, super-resolution, and inpainting tasks. Our main goal is to benchmark against DIP as the foundational implicit prior, while also comparing to recent strong unsupervised and task-specific SOTA methods. DSP is proposed with the goal of not to beat all SOTA unsupervised techniques but to introduce a new direction for frequency-domain implicit models. We report PSNR and visual results (with zoom-in views) on standard benchmarks (mairal2009non, ; zeyde2010single, ; huang2015single, ; ma2022regionwise, ; yu2019free, ), using the same iteration schedule for DIP and DSP. All experiments run on a Tesla T4 GPU (15GB RAM).
Blind Denoising. In the blind denoising task, the forward measurement operator is the identity operator. We compare the denoising performance (measured in PSNR, dB) of our proposed DSP method with several established unsupervised approaches, including DIP ulyanov2018deep , CBM3D lebrun2012analysis , Noise2Self batson2019noise2self , Wavelet-based shrinkage chang2000adaptive , Total Variation (TV) denoising chambolle2004algorithm ; getreuer2012rudin , WIRE saragadam2023wire , and DS-N2N bai2025dual , across a variety of test images. As shown in Table 1, DSP achieves notable improvements over the baselines, particularly on fine-textured or complex regions such as “Plane” and “Brick Wall,” with gains of +1.69 dB and +1.53 dB over DIP, respectively. This suggests that DIP tends to overfit to noise or overly smooth detailed textures, whereas DSP better preserves structural details. Visual comparisons in Figure 1 further confirm these findings: DSP delivers cleaner reconstructions with enhanced texture fidelity, avoiding the patchy artefacts commonly observed in Noise2Self and the grid-like behaviour and over-smoothing typical of CBM3D in highly structured regions.
Image Restoration. In the image restoration task, a Bernoulli mask is applied as the forward measurement operator.
We compare our DSP method against baselines including DIP, and Convolutional Sparse Coding (CSC) wohlberg2016boundary implemented with wohlberg2016sparse ; wohlberg2017sporco . As shown in Table 2, DSP consistently outperforms these methods. For example, on the “Barbara” image, which contains fine patterns and directional textures, DSP yields a higher PSNR than both DIP and TV, indicating better preservation of structured detail. Qualitative comparisons in Figure 2 further reveal that DSP reconstructs finer textures and edges more faithfully, whereas TV introduce artefacts in regions, CSC tend to produce over-smoothed results and DIP kept masked artifact with non smoothed edge features. These results highlight the capability of DSP to recover complex structures even under aggressive masking.
@@ 3 Experimental Results
Wavelet chang2000adaptive 2000 21.37 26.54 24.76 22.19 22.16 28.91 28.45
TV chambolle2004algorithm 2004 23.90 23.55 23.90 23.40 23.29 25.04 28.70
CBM3D lebrun2012analysis 2012 19.73 21.95 20.57 19.86 20.77 26.28 27.29
TV Bregman getreuer2012rudin 2012 17.45 22.70 21.29 18.58 19.53 23.26 24.28
DIP ulyanov2018deep 2018 24.26 27.65 24.27 22.95 22.77 30.99 28.83
Noise2Self batson2019noise2self 2019 22.81 25.19 23.81 21.94 21.31 28.92 28.64
WIRE saragadam2023wire 2023 19.92 25.41 22.93 20.35 21.57 29.22 29.64
DS-N2N bai2025dual 2025 25.45 27.96 25.77 23.56 23.36 30.65 30.40
@@
CSC wohlberg2016boundary 2016 28.22 33.68 28.22 28.88 28.02 31.64 29.18
DIP ulyanov2018deep 2018 29.61 37.22 32.68 32.55 29.49 31.83 30.97
@@
Table 3: PSNR (dB) comparison across super-resolution methods (Bicubic, DIP, TV, LapSRN, ResShift, SinSR (with single step and 15 steps in wang2024sinsr ), and ASID with our proposed DSP on various test images. Best values for both supervised and unsupervised groups are highlighted in bold.
       Method Zebra PPT Flower Comic Bird
×$\times 4$$\times 8$$\times 4$$\times 8$$\times 4$$\times 8$$\times 4$$\times 8$$\times 4$$\times 8$
Unsup Bicubic 23.10 18.55 20.31 17.17 23.76 20.52 20.20 17.77 28.44 23.30
TV rudin1992nonlinear 1992 23.51 18.68 20.51 17.32 23.99 20.66 20.32 17.84 28.75 23.67
DIP ulyanov2018deep 2018 23.83 19.55 22.47 18.72 24.25 21.31 20.75 18.44 29.13 24.94
DSP (ours) 2025 24.31 19.68 22.99 18.84 24.57 21.41 20.90 18.48 30.27 25.04
@@
ResShift yue2023resshift 2023 25.03 23.43 23.52 20.86 30.00
SinSR wang2024sinsr 2024 21.51 22.22 22.43 18.62 26.55
SinSR 15 2024 21.67 22.10 22.28 19.09 26.75
ASID park2025efficient 2025 18.90 22.59 23.48 18.90 32.59
Table 4: Performance comparison across supervised methods (GLCIC and Prefpaint) and unsupervised methods (DIP and DSP (ours)) on inpainting tasks. Best values per image are highlighted.
  Method Vase Street Face Beach Shore Glider Backstreet Library
Unsup DIP ulyanov2018deep 2018 28.32 22.18 22.91 19.89 29.43 29.46 23.85 19.06
DSP (ours) 2025 29.44 22.63 23.52 20.06 31.27 29.96 26.34 19.37
[…]
Prefpaint liu2024prefpaint 2024 22.74 23.21 25.43 19.85 27.68 27.40 23.04 16.80
Figure 3: Comparison of super resolution task with 4$\times$ upscale on the “Zebra” image among Bicubic, TV, DIP, and the supervised methods (LapSRN, ResShift, and SinSR) with our proposed DSP.
@@
Super Resolution. For the super resolution task, the forward measurement operator corresponds to downsampling with fixed scaling factors. We compare our proposed DSP method with a range of both unsupervised and supervised baselines. The unsupervised group includes DIP ulyanov2018deep , Total Variation (TV) rudin1992nonlinear , and Bicubic upsampling, while the supervised group comprises LapSRNtai2017image , ResShift yue2023resshift , SinSR wang2024sinsr and ASID park2025efficient . Since supervised methods are trained for fixed resolutions, only the officially available ×$\times 4$ upscaling is evaluated. As supervised methods are trained for fixed resolutions, only the officially available ×$\times 4$ results are reported. In contrast, DSP operates at arbitrary scales without any task-specific pretraining. As shown in Table 3, DSP consistently achieves the highest PSNR across all test images in both ×$\times 4$ and ×$\times 8$ settings among unsupervised methods. Notably, on the “Bird” image, DSP attains 30.27dB, outperforming DIP by over 1dB and even surpassing some supervised baselines such as SinSR and ResShift. Figure 3 offers a visual comparison on the “Zebra” image at ×$\times 8$ upscaling. While supervised methods tend to produce smoother outputs, they often oversimplify textures due to reliance on generic learned priors. In contrast, DSP better preserves fine structures such as the zebra’s stripes and background foliage, showing sharper detail and fewer artefacts than DIP. These results highlight DSP’s ability to recover image-specific features and fine textures in a single-shot, unsupervised setting—making it a flexible and robust approach for real-world super resolution tasks.
Inpainting. We evaluate DSP on the inpainting task using a range of unsupervised and supervised baselines. The unsupervised methods include DIP ulyanov2018deep , while the supervised methods are UNet rudin1992nonlinear , ResNet ronneberger2015u ; he2016deep , GLCIC iizuka2017globally and Prefpaint liu2024prefpaint . As shown in Table 4, DSP outperforms DIP across all test images, with consistent PSNR improvements. On several images such as “Shore” and “Library”, DSP even surpasses supervised methods, which confirm that DSP effectively leverages image-specific structures without requiring external training data. Qualitative results on the “Library” image are shown in Figure 4. DIP and ResNet produce dark twisted regions where the mask is present, indicating unstable reconstructions in under-constrained areas. UNet offers smoother outputs but fails to recover semantically plausible structures. GLCIC, a GAN-based method, introduces edge artefacts near mask boundaries, likely due to adversarial limitations. Although Prefpaint yields clean and realistic completions, it diverges from the ground truth content, aligning instead with learned aesthetic preferences. In contrast, DSP generates visually coherent and structurally faithful reconstructions, striking a strong balance between realism and fidelity, despite operating entirely without supervision or pretrained models.
@@
Early Stopping Analysis. We analyse the iterative progression of DIP and DSP on the “Plane” image from the blind denoising task, as visualised in Figure 5. DIP reaches its peak PSNR around 1800 iterations but then quickly dropped to 28.83 dB by iteration 3000 due to overfitting. In contrast, DSP improves steadily throughout optimisation, reaching 30.52 dB at 3000 iterations. The visual comparison shows that DSP produces increasingly cleaner and structurally consistent reconstructions, while DIP begins to introduce artefacts after its peak. This highlights DSP’s stability and robustness across iterations, especially in the absence of manual early stopping.

### v2

@@ 2.3 The Spectral Stability Law: Robust Early-Stopping Dynamics of DSP
Wavelet [chang2000adaptive] 2000 21.37 26.54 24.76 22.19 22.16 28.91 28.45
TV [chambolle2004algorithm] 2004 23.90 23.55 23.90 23.40 23.29 25.04 28.70
CBM3D [lebrun2012analysis] 2012 19.73 21.95 20.57 19.86 20.77 26.28 27.29
TV Bregman [getreuer2012rudin] 2012 17.45 22.70 21.29 18.58 19.53 23.26 24.28
DIP [ulyanov2018deep] 2018 24.26 27.65 24.27 22.95 22.77 30.99 28.83
Noise2Self [batson2019noise2self] 2019 22.81 25.19 23.81 21.94 21.31 28.92 28.64
WIRE [saragadam2023wire] 2023 19.92 25.41 22.93 20.35 21.57 29.22 29.64
DS-N2N [bai2025dual] 2025 25.45 27.96 25.77 23.56 23.36 30.65 30.40
@@
CSC [wohlberg2016boundary] 2016 28.22 33.68 28.22 28.88 28.02 31.64 29.18
DIP [ulyanov2018deep] 2018 29.61 37.22 32.68 32.55 29.49 31.83 30.97
@@
Blind Denoising. In the blind denoising task, the forward measurement operator is the identity operator. We compare the denoising performance (measured in PSNR, dB) of our proposed DSP method with several established unsupervised approaches, including DIP [ulyanov2018deep], CBM3D [lebrun2012analysis], Noise2Self [batson2019noise2self], Wavelet-based shrinkage [chang2000adaptive], Total Variation (TV) denoising [chambolle2004algorithm, getreuer2012rudin], WIRE [saragadam2023wire], and DS-N2N [bai2025dual], across a variety of test images. As shown in Table 2.3, DSP achieves notable improvements over the baselines, particularly on fine-textured or complex regions such as “Plane” and “Brick Wall,” with gains of +1.69 dB and +1.53 dB over DIP, respectively. This suggests that DIP tends to overfit to noise or overly smooth detailed textures, whereas DSP better preserves structural details. Visual comparisons in Figure 2 further confirm these findings: DSP delivers cleaner reconstructions with enhanced texture fidelity, avoiding the patchy artefacts commonly observed in Noise2Self and the grid-like behaviour and over-smoothing typical of CBM3D in highly structured regions. Additionally, Figure 3 illustrates CBM3D and Noise2Self produce visible grid-like artifacts, with Noise2Self further exhibiting noticeable over-smoothing, whereas DIP allows residual noise to persist—likely due to the iterative process gradually pulling the reconstruction closer to the noisy input—while DSP successfully preserves the sharp structure and fine details.
Image Restoration. In the image restoration task, a Bernoulli mask is applied as the forward measurement operator. We compare our DSP method against baselines including DIP, and Convolutional Sparse Coding (CSC) [wohlberg2016boundary] implemented with [wohlberg2016sparse, wohlberg2017sporco]. As shown in Table 2.3, DSP consistently outperforms these methods. For example, on the “Barbara” image, which contains fine patterns and directional textures, DSP yields a higher PSNR than both DIP and TV, indicating better preservation of structured detail. Qualitative comparisons in Figure 4 further reveal that DSP reconstructs finer textures and edges more faithfully, whereas TV introduce artefacts in regions, CSC tend to produce over-smoothed results and DIP kept masked artifact with non smoothed edge features. These results highlight the capability of DSP to recover complex structures even under aggressive masking.
@@
Super Resolution. For the super resolution task, the forward measurement operator corresponds to downsampling with fixed scaling factors. We compare our proposed DSP method with a range of both unsupervised and supervised baselines. The unsupervised group includes DIP [ulyanov2018deep], Total Variation (TV) [rudin1992nonlinear], and Bicubic upsampling, while the supervised group comprises LapSRN[tai2017image], ResShift [yue2023resshift], SinSR [wang2024sinsr] and ASID [park2025efficient]. Since supervised methods are trained for fixed resolutions, only the officially available $\times 4$ upscaling is evaluated. As supervised methods are trained for fixed resolutions, only the officially available $\times 4$ results are reported. In contrast, DSP operates at arbitrary scales without any task-specific pretraining. As shown in Table LABEL:psnr_comparison_sr, DSP consistently achieves the highest PSNR across all test images in both $\times 4$ and $\times 8$ settings among unsupervised methods. Notably, on the “Bird” image, DSP attains 30.27dB, outperforming DIP by over 1dB and even surpassing some supervised baselines such as SinSR and ResShift. Figure 2.3 offers a visual comparison on the “Zebra” image at $\times 8$ upscaling. While supervised methods tend to produce smoother outputs, they often oversimplify textures due to reliance on generic learned priors. In contrast, DSP better preserves fine structures such as the zebra’s stripes and background foliage, showing sharper detail and fewer artefacts than DIP. These results highlight DSP’s ability to recover image-specific features and fine textures in a single-shot, unsupervised setting—making it a flexible and robust approach for real-world super resolution tasks.
Inpainting. We evaluate DSP on the inpainting task using a range of unsupervised and supervised baselines. The unsupervised methods include DIP [ulyanov2018deep], while the supervised methods are UNet [rudin1992nonlinear], ResNet [ronneberger2015u, he2016deep], GLCIC [iizuka2017globally] and Prefpaint [liu2024prefpaint]. As shown in Table LABEL:inpaint_psnr_comparison, DSP outperforms DIP across all test images, with consistent PSNR improvements. On several images such as “Shore” and “Library”, DSP even surpasses supervised methods, which confirm that DSP effectively leverages image-specific structures without requiring external training data. Qualitative results on the “Library” image are shown in Figure 2.3. DIP and ResNet produce dark twisted regions where the mask is present, indicating unstable reconstructions in under-constrained areas. UNet offers smoother outputs but fails to recover semantically plausible structures. GLCIC, a GAN-based method, introduces edge artefacts near mask boundaries, likely due to adversarial limitations. Although Prefpaint yields clean and realistic completions, it diverges from the ground truth content, aligning instead with learned aesthetic preferences. In contrast, DSP generates visually coherent and structurally faithful reconstructions, striking a strong balance between realism and fidelity, despite operating entirely without supervision or pretrained models.
Early Stopping Analysis. We analyse the iterative progression of DIP and DSP on the “Plane” image from the blind denoising task, as visualised in Figure 2.3. DIP reaches its peak PSNR around 1800 iterations but then quickly dropped to 28.83 dB by iteration 3000 due to overfitting. In contrast, DSP improves steadily throughout optimisation, reaching 30.52 dB at 3000 iterations. The visual comparison shows that DSP produces increasingly cleaner and structurally consistent reconstructions, while DIP begins to introduce artefacts after its peak. This highlights DSP’s stability and robustness across iterations, especially in the absence of manual early stopping.